You are given the following data.

S(s) + 3/2 O2(g) → SO3(g) ΔH = -395.2 kJ

2 SO2(g) + O2(g) → 2 SO3(g) ΔH = -198.2 kJ

Calculate ΔH for the reaction below.

S(s) + O2(g) → SO2(g)
1
Expert's answer
2020-11-02T08:52:09-0500

S(s) + 3/2O2(g) → SO3(g) (I) ΔH1 = -395.2 kJ

2SO2(g) + O2(g) → 2SO3(g) (II) ΔH2 = -198.2 kJ

(I) – 1/2(II)

S(s) + 3/2O2(g) + SO3 → SO2(g) + 1/2O2(g) + SO3

S(s) + O2(g) → SO2(g)

ΔH = ΔH1 – 1/2ΔH2

ΔH = -395.2 +1/2(198.2) = -296.1 kJ